OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

1- I think its this one 4 5 25

OpenStudy (anonymous):

All you have to do is distribute 3(x+1)=-2(x-1)+6 3x+3=-2x+2+6 Add Like terms 3x+3=-2x+8 add 2x on both sides +2x +2x ________________ 5x+3=8 Subtract 1 on both sides -3 -3 ___________________ 5x=5 Now just divide 5 on both sides x=?

OpenStudy (anonymous):

So it would be 1, right?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Thanks a million

OpenStudy (anonymous):

No Problem!